# PBEBench: A Multi-Step Programming by Examples Reasoning Benchmark inspired by Historical Linguistics

### v3

@@ Abstract
Although many benchmarks evaluate the reasoning abilities of Large Language Models (LLMs) within domains such as mathematics, coding, or data wrangling, few abstract away from domain specifics to examine reasoning as a capability in and of itself. We contribute a novel type of benchmark evaluating the inductive reasoning capabilities of LLMs that is inspired by the forward reconstruction task from historical linguistics but is formulated in an extremely simple, general way (in the form of Programming by Examples or PBE). The task involves generating a cascade of simple string rewrite programs to transform a given list of input strings into a list of desired output strings. We present a fully automated pipeline that programmatically generates problems of this type with controllable difficulty (varying, for example, ground truth cascade lengths), enabling scalable evaluation of reasoning models while avoiding contamination. Using this approach, we construct two benchmarks: PBEBench-Lite, which efficiently stratifies models of varying capabilities, and PBEBench, which requires models to induce programs more similar (in complexity) to those constructed by historical linguists. Our experiments reveal a substantial performance gap between models that leverage test-time compute or LCoT (long chain-of-thought) reasoning and those that do not. Moreover, although recent models such as gpt-oss-120b and GPT-5 show promise, the solve rate for both of them drops below 5% for hard instances of the PBEBench dataset (ground truth cascade lengths of 20 and 30, respectively), falling well short of realistic historical linguistics requirements even with computationally expensive, popular scaling techniques from the PBE and reasoning literature. Additionally, we also study the effectiveness of different scaling strategies and the impact of various hyperparameters on the difficulty of the generated data using gpt-oss-120b, the best-performing open-source model. We plan to open-source our code and benchmark snapshots to enable reproducibility and future work.
@@ 1 Introduction
You are trying to refactor a large codebase, implementing changes like replacing all instances of foo with bar, bar with baz, and baz with foo. When you tell your senior colleague, a seasoned Unix hacker, what you are doing, he smiles and says that he can (after inspecting the originals and your examples of refactored code) write a sed (stream editor) script that will update the entire codebase. His task involves not only constructing the appropriate regular expression replacements, but also placing them in the correct sequence. For example, s/bar/baz/ (replace all instances of bar with baz) has to be ordered before s/foo/bar/ (otherwise, all instances that were originally foo would become baz). Likewise, there must be a substitution like s/baz/TEMP/ that occurs at the beginning of the the “cascade” and one like s/TEMP/foo that occurs at the end (yielding a circular mapping). At a deep level, this problem is the same as one from historical linguistics: given a set of pairs consisting of a word in an ancestral language and a word in a modern language, define a sequence of local string transformations that convert the ancestral words to the modern words (a task called forward reconstruction). In both cases, one is given a collection of inputs $\vec{i}$ (input files or ancestral words) and a collection of outputs $\vec{o}$ to which the inputs should be mapped by a program $\vec{p}$ consisting of a sequence of incremental transformations. Both are clear instances of programming example, Naik et al. (2025b) (the inspiration for the current work) having made this explicit for forward reconstruction.
An important observation of Naik et al. (2025b) is that this class of tasks is challenging for LLMs. Because of this, and because it can be reduced to the barest formal machinery, it provides an ideal framework for comparing the ability of language models to construct (by induction) sequential plans. The principle contribution of the paper is a benchmark of this type. It has a number of desirable properties underrepresented in current benchmarks for reasoning: (1) Unlike most reasoning benchmarks (see 2), it is not bound to a specific domain and does not benefit from non-trivial domain knowledge. (2) Its complexity and difficulty emerge from a very simple framework with provable logic. (3) This emergent complexity includes four fundamental challenges (to capture the relationships Kiparsky (1968; 1971) called feeding, bleeding, counter-feeding, and counter-bleeding) which can be manipulated and quantified. (4) The benchmark is immune to data contamination and resistant to saturation because new test instances can be generated, with controlled difficulty very efficiently. (5) It compares the ability of models to perform a class of practically and scientifically interesting tasks rather than focusing on comparisons between model and human abilities.
Our study addresses the following questions:
RQ1: Which models are good at inductive reasoning? We benchmark 13 reasoning and 8 strong non-reasoning models on PBEBench-Lite (a lightweight version of the proposed benchmark) to determine whether long chain of thought (LCoT) reasoning is beneficial in this task.
RQ2: What makes a PBE instance hard? We use logistic regression and factorial analysis to determine how much factors like ground truth cascade length, types of relations between rewrites, etc., impact the solution difficulty of dataset problems, based on solution success across all models. We also create a snapshot similar to PBEBench-Lite but with 50 examples per PBE instance (PBEBench-Lite-MoreEg) to evaluate the impact of having more examples on solution success rate.
RQ3: Could the strongest LLMs solve problems requiring long programs (like real-world forward reconstruction problems)? We pick the strongest open and closed-source models from the PBEBench-Lite results and analyze them on PBEBench (the full benchmark).
RQ4: What scaling strategies work best for solution search? We investigate which of the two solution search scaling strategies discussed in Section 3.2 yields more “bang for the buck” by holding the sampling budget constant and varying the maximum sequence length and vice-versa.
@@ 2 Related Work
Programming By Example. Programming by Example (PBE) (Gulwani, 2010) is a well-known and intuitive paradigm in program synthesis research, where programs are inferred solely from a small set of input-output pairs. Early symbolic approaches relied on domain-specific languages (DSLs) and constraint solving: FlashFill uses a string-transformation DSL to automate spreadsheet tasks (Gulwani, 2011), while Syntax-Guided Synthesis (SyGuS) constrains program search via grammar specifications (Alur et al., 2013). Later, DeepCoder (Balog et al., 2017) guided search using learned function predictions, and RobustFill (Devlin et al., 2017) trained sequence-to-sequence models to directly emit DSL programs. More recently, Large Language Models have shown few-shot capabilities in code generation (Chen et al., 2021a; Guo et al., 2024) and test case generation (Li & Yuan, 2024), but on traditional PBE tasks they often struggle with out-of-distribution examples, and improve only after fine-tuning on the target distribution (Li & Ellis, 2024; Naik et al., 2025b).
Inducing Context-Sensitive Grammars in LLMs. Attempts to induce string-rewrite rules from data have a long history (Gildea & Jurafsky, 1995); discussions of the formal properties of these rules in linguistics go back farther, including (pivotal to this study) the discovery of feeding and bleeding relationships between such rules Kiparsky (1968; 1971). More recently, formal language benchmarks have shown that RNNs can outperform transformers on formal language tasks involving certain classes of grammars (Butoi et al., 2025). Morphophonological probes (Borenstein, 2024) reveal that while models fit the training data, they often default to heuristics, highlighting limitations in true rule induction. Recent work analyzing LLM-based hypothesis search further documents systematic errors in rule induction, showing that models often produce oversimplified hypotheses that fail to generalize (Parab et al., 2025). Naik et al. (2024) demonstrate that LLMs can induce sound laws, generalizing the method to full context-sensitive program synthesis (Naik et al., 2025b). However, none of these prior works have proposed provably correct algorithms for detecting feeding and bleeding relations (a unique contribution of this work; see Appendix C).
Reasoning and Induction Benchmarks. A range of benchmarks test reasoning and PBE skills. Code-centric suites include HumanEval and MBPP (Chen et al., 2021a; Austin et al., 2021), while PBE-style tasks appear in FlashFill datasets (Gulwani, 2011). Newer benchmarks include TESTGENEVAL (Jain et al., 2025), that targets unit tests generation. CODE2BENCH (Zhang et al., 2025) has been proposed to improve scalability in benchmark generation for software programs, while Case2Code (Shao et al., 2025) introduces a scalable synthetic dataset for program synthesis (requiring models to infer code from input-output examples). However, these approaches are tied to software engineering, where datasets demand both domain knowledge and reasoning, introducing unequal prior exposure across models and preventing contamination-free evaluation of the knowledge-free reasoning ability we aim to study. Linguistic and reasoning benchmarks such as HotpotQA (Yang et al., 2018), DROP (Dua et al., 2019), and GSM8K (Cobbe et al., 2021) stress multi-step and mathematical reasoning. System-2 reasoning is evaluated in BIG-Bench Hard (Suzgun et al., 2022), where chain-of-thought prompting improves performance significantly. Compositional generalization splits like SCAN (Lake & Baroni, 2018) and CFQ (Keysers et al., 2020) reveal extrapolation gaps. Visual PBE in ARC (Chollet, 2019) is a non-language analog to our task. However, our benchmark is significantly more resistant to leakage, easily scalable, and offers a mechanism to modulate task difficulty by simple adjustment of the generation parameters. To our knowledge, no benchmark targets multi-step synthesis of string-rewriting programs; PBEBench fills this niche by unifying PBE and compositional generalization in a single benchmark. The Illusion of Thinking (Shojaee et al., 2025) highlights reasoning collapse and out-of-token issues in complex tasks, with subsequent work (Lawsen, 2025) noting methodological limitations. We encounter similar issues during evaluation, but counter them by scaling chain-of-thought budgets, adding additional sampling when needed, and extending timeouts with parallelized inference for efficiency.
@@ 3 Methodology
Our approach consists of two components: (1) a problem proposer and (2) a problem solver. The proposer is a programmatic system that generates PBE instances of controllable difficulty, forming the core of our dynamic benchmarking framework. It scalably and reliably produces novel, contamination-free data of the desired complexity. The solver is any system under evaluation. In this work, we focus on state-of-the-art open and closed-source reasoning LLMs to examine the limits of their inductive reasoning capabilities and to assess whether LLM-based methods can be deployed for realistic historical linguistics sound law induction tasks.
Figure 1: Structure of the $\vec{i},\vec{p},\vec{o}$> triples dataset. The relations between the programs in the sequence are shown by the BFCC DAG, and the simplified BF DAG replaces the counterfactual relations with a reversed simple link.
@@ 3.1 Problem Proposer
We construct our benchmark programmatically by sampling inputs from a distribution of strings $\mathcal{I}$ and a sequence of string rewrite programs from a distribution $\mathcal{P}$. Each string rewrite program acts like a find and replace function and substitutes all occurrences of substring $\alpha$ with a substring $\beta$ (details in Section 3.1.2). We apply the sampled cascade of string rewrite programs $\vec{p}=\langle p_{1},\dots,p_{m}\rangle\in\mathcal{P}$ to the sampled inputs $\vec{i}=\langle i_{1},\dots,i_{n}\rangle\in\mathcal{I}$, producing outputs $\vec{o}=\{p_{m}(\dots p_{1}(i_{1})),\dots,p_{m}(\dots p_{1}(i_{n}))\}$. For convenience, we write this transformation as $\vec{o}=\vec{p}(\vec{i})$. Each instance in our dataset is defined by the triplet $\langle\vec{i},\vec{p},\vec{o}\rangle$, consisting of the inputs, the program cascade, and the corresponding outputs. Additionally, we record metadata such as the interactions between programs (e.g., bleeding, feeding, etc.).
Our data generation algorithm is parameterized by: the number of examples $n$ ($|\vec{i}|=|\vec{o}|=n$); the input alphabet $\Sigma$ used to generate both the inputs and the string rewrite programs; the minimum and maximum input string lengths $l_{min}$ and $l_{max}$; the minimum and maximum cascade lengths $L_{min}$ and $L_{max}$ ($L_{min}\leq|\vec{p}|\leq L_{max}$); the minimum and maximum length of strings $s_{min}$ and $s_{max}$ ($s_{min}\leq|\alpha|,|\beta|\leq s_{max}$) used by the string rewrite programs and the desired dataset size $D$. We denote a benchmark snapshot generated with specific parameter values as $\mathcal{D}(n,\Sigma,L_{min},L_{max},l_{min},l_{max},s_{min},s_{max},D)$.
@@ 3.1.1 Input Sampling
Our input sampling procedure is parameterized by $(n,\Sigma,l_{min},l_{max})$ as defined in Section 3.1. We build the input sequence $\vec{i}=\langle i_{1},\dots,i_{n}\rangle$ by independently sampling each input $i_{k}$ for $1\leq k\leq n$. For each $i_{k}$, we first sample its length from a uniform discrete distribution, $|i_{k}|\sim\mathrm{Unif}\{l_{min},\dots,l_{max}\}$, where $|i_{k}|$ denotes the length of the string $i_{k}$. We then generate the string itself by sampling $|i_{k}|$ characters independently with replacement from the alphabet $\Sigma$, i.e., $i_{k}\sim\mathrm{Unif}(\Sigma)^{|i_{k}|}$.
@@ 3.1.2 Program Sampling and Output Generation
Program sampling is parameterized by $(\Sigma,L_{min},L_{max},s_{min},s_{max})$. We begin by programmatically selecting a cascade length $L$ such that $L_{min}\leq L\leq L_{max}$. Next, we construct a sequence of $L$ string rewrite programs $\vec{p}=\langle p_{1},\dots,p_{L}\rangle$, where each program is of the form $p_{k}=\mathrm{replace}(\alpha_{k},\beta_{k})$. Here, $\mathrm{replace}$ has the same semantics as Python’s built-in replace() method for strings: the substring $\alpha_{k}$ is replaced by $\beta_{k}$, with the restriction that $\alpha_{k}$ is non-empty. To sample each program $p_{k}$, we generate $\alpha_{k}$ and $\beta_{k}$ independently, following a procedure analogous to input sampling:
@@
    Sample $\alpha_{k}$ uniformly from the set of substrings of length $|\alpha_{k}|$ in the intermediate input vector $\vec{i}_{k-1}=p_{k-1}(\dots p_{1}(\vec{i}))$ (where $\mathrm{Substr}_{s}(\vec{i}_{k-1})$ denotes the set of all substrings of length $s$ in $\vec{i}_{k-1}$): $\alpha_{k}\sim\mathrm{Unif}(\mathrm{Substr}_{|\alpha_{k}|}(\vec{i}_{k-1}))$ (note that $\vec{i}_{0}=\vec{i}$).
@@
The outputs are then generated by executing the program cascade $\vec{p}$ over the inputs $\vec{i}$ as $\vec{o}=\vec{p}(\vec{i})$

### v4

@@ Abstract
While many benchmarks evaluate the reasoning abilities of Large Language Models (LLMs), few isolate reasoning as a capability independent of domain knowledge. We introduce a new benchmark for inductive reasoning inspired by Sound Law Induction (SLI) in historical linguistics and formulated in a simple multi-step Programming by Example (PBE) framework. The task requires inducing a cascade of string rewrite programs that transform inputs into target outputs. We present PBEBench, a fully automated evaluation approach that generates such problems with controllable difficulty and ordering constraints, enabling scalable and contamination-resistant evaluation of sequential inductive reasoning. Using this approach, we construct three datasets that show a large gap between models that leverage test-time compute or long chain-of-thought reasoning and those that do not. Although recent models such as GPT-5 and gpt-oss-120b show promise, solve rates remain below 5% on hard PBEBench instances with long program cascades, even under computationally expensive scaling strategies. Finally, we show that PBEBench scores are more predictive of performance on real SLI than are other inductive reasoning benchmarks. We will release code and data to support further research.
PBEBench: A Multi-Step Programming by Examples Reasoning Benchmark inspired by Historical Linguistics
Atharva Naik1, Darsh Agrawal1, Prakam3, Yash Mathur3, Manav Kapadnis1, Yuwei An1, Clayton Marr2, Carolyn Rose1, David Mortensen1 1Carnegie Mellon University, 2Ohio State University, 3Independent Researcher {arnaik,darsha,mkapadni,yuweia,cprose,dmortens}@cs.cmu.edu
@@ 1 Introduction
[…]
Naik et al. (2025b, 2024) show that LLMs struggle with SLI, but benefit more from training on logically consistent synthetic sound laws than on real sound laws from Index Diachronica. However, these works do not explain why this is the case. We hypothesize that the difficulty arises not from individual rewrite operations, but from inducing sequences of interacting transformations whose correctness depends on precise ordering and unobserved intermediate states, explaining the advantage of structurally consistent synthetic data. Despite minimal formal machinery, the problem exhibits rich interaction structure, making it a natural testbed for multi-step inductive reasoning and sequential planning in a knowledge-free setting Ma et al. (2024). We propose PBEBench, an automated approach towards generating such data automatically and scalably. PBEBench exhibits several desirable properties missing in existing reasoning benchmarks: (1) it does not rely on specialized background knowledge (see Section 2); (2) its difficulty emerges from a simple, formally specified, finite, problem space of sequential string rewrites; (3) it supports explicit control of fundamental interaction patterns—feeding, bleeding, counter-feeding, and counter-bleeding—studied in historical phonology Kiparsky (1968, 1971) (termed as BFCC relations here); (4) it is inherently resistant to data contamination and saturation, as new instances with controlled difficulty can be generated scalably; and (5) it evaluates models on practically and scientifically meaningful induction problems rather than human–model comparisons.
Our study investigates the following hypotheses:
H1: LLMs struggle with multi-step inductive reasoning. We benchmark 13 reasoning and 8 non-reasoning models on three PBEBench datasets to study reasoning bottlenecks in LLMs. The knowledge-free setting suggests that reasoning models should outperform non-reasoning ones.
H2: Difficulty in our dataset is driven by interaction structure. We hypothesize that while inducing a single string rewrite is straightforward, LLMs struggle to compose long cascades, especially under complex BFCC ordering constraints. We test this by contrasting increased per-step complexity (via more examples) with longer cascades and richer BFCC interactions.
H3: Performance on PBEBench predicts performance on real SLI. We evaluate all open-source LLMs on real SLI data and contemporary synthetic inductive reasoning benchmarks, and compare correlations to identify which benchmarks best predict real SLI performance.
H4: Scaling the thinking budget is more effective than increasing the sampling budget. We compare two scaling strategies: increasing the thinking budget (maximum sequence length) and the sampling budget, as described in Section 3.2, to determine which yields more “bang for the buck” under a fixed compute budget.
@@ 2 Related Work
Programming By Example. Programming by Example (PBE) (Gulwani, 2010) is a program synthesis paradigm where programs are inferred from input–output pairs. Early symbolic approaches rely on domain-specific languages and constraints: FlashFill uses a string-transformation DSL (Gulwani, 2011), while Syntax-Guided Synthesis (SyGuS) restricts program search via grammars (Alur et al., 2013). DeepCoder (Balog et al., 2017) learns function predictions to guide search, and RobustFill (Devlin et al., 2017) trains sequence-to-sequence models to directly emit DSL programs. More recently, Large Language Models show few-shot capability in code generation (Chen et al., 2021a; Guo et al., 2024) and test case generation (Li and Yuan, 2024), but on traditional PBE, they struggle with out-of-distribution data and typically improve after fine-tuning on the target distribution (Li and Ellis, 2024; Naik et al., 2025b).
Inducing Context-Sensitive Grammars in LLMs. Attempts to induce string-rewrite rules from data have a long history (Gildea and Jurafsky, 1995), with linguistic analyses of their formal properties dating back further Kiparsky (1968, 1971). Naik et al. (2024) show that LLMs can induce sound laws, extending this to full context-sensitive program synthesis (Naik et al., 2025b). However, these works do not provide provably correct algorithms for detecting feeding and bleeding relations, which is a unique contribution of this work (see Appendix B).
Reasoning and Induction Benchmarks. Several benchmarks have been proposed to evaluate reasoning and PBE capabilities. Code-centric suites such as HumanEval and MBPP assess function generation (Chen et al., 2021a; Austin et al., 2021), while FlashFill-style datasets target example-driven string transformations (Gulwani, 2011). More recent benchmarks emphasize software-engineering workflows (Jain et al., 2025; Zhang et al., 2025; Shao et al., 2025). However, software engineering tasks often entangle general reasoning with domain-specific knowledge (of tools and libraries), leading to uneven prior exposure and complicating contamination-free evaluation. Beyond code, multi-step, system-2, and mathematical reasoning are explored by HotpotQA, DROP, BIG-Bench Hard, and GSM8K (Yang et al., 2018; Dua et al., 2019; Suzgun et al., 2022; Cobbe et al., 2021). Inductive rule learning and compositional generalization are probed via CLUTRR and SLR (Sinha et al., 2019; Helff et al., 2025), bAbI and KOR-Bench (Weston et al., 2015; Ma et al., 2024), and extrapolative splits such as SCAN and CFQ (Lake and Baroni, 2018; Keysers et al., 2020); ARC provides a non-language analog (Chollet, 2019). In contrast, PBEBench targets multi-step synthesis of string-rewriting cascades, is easily scalable and leakage-resistant, and supports difficulty control through simple generation parameters.
@@ 3 Methodology
Our approach comprises two components: (1) a problem proposer and (2) a problem solver (Fig. 1(a)). The proposer is a symbolic system that generates PBE instances with controllable difficulty and forms the core of our dynamic benchmarking framework. It scalably produces novel, contamination-free input, output, and program triples of controllable difficulty. The solver corresponds to any system under evaluation. We benchmark state-of-the-art LLMs on inductive reasoning using a program reordering, multi-step PBE task, and additionally evaluate them on human-curated, low-resource Sound Law Induction (SLI) data.
(a) Proposer–solver architecture
(b) Structure of a PBE data instance
Figure 1: Overview of PBEBench. (a) A symbolic proposer generates PBE instances of controllable difficulty, and a solver attempts to recover the underlying program sequence or ordering. (b) Each instance consists of $\vec{\imath},\vec{p},\vec{o}$> triples, with program relations encoded by the BFCC DAG and a simplified BF variant that replaces counterfactual relations with a reversed link.
@@ 3.1 Problem Proposer
The symbolic proposer (Fig 1(a), top) constructs data for the multi-step PBE task by sampling inputs from a distribution of strings $\mathcal{I}$ and a sequence of string rewrite programs from a distribution $\mathcal{P}$. Each program acts as a find-and-replace function, substituting all occurrences of a substring $\alpha$ with a substring $\beta$ (details in Section 3.1.2). The proposer applies the sampled cascade of programs $\vec{p}=\langle p_{1},\dots,p_{m}\rangle\in\mathcal{P}$ to the sampled inputs $\vec{\imath}=\langle i_{1},\dots,i_{n}\rangle\in\mathcal{I}$, producing outputs $\vec{o}=\langle p_{m}(\dots p_{1}(i_{1})),\dots,p_{m}(\dots p_{1}(i_{n}))\rangle$. For convenience, we write this transformation as $\vec{o}=\vec{p}(\vec{\imath})$. Each dataset instance (Fig 1(a), second from top) is defined by the triplet $\langle\vec{\imath},\vec{p},\vec{o}\rangle$, consisting of the inputs, the program cascade, and the corresponding outputs. In addition, the proposer records metadata describing interactions between programs (e.g., bleeding, feeding), illustrated in Fig 1(b).
The proposer is parameterized by the number of examples $n$ ($|\vec{\imath}|=|\vec{o}|=n$; $n=5$ in Fig 1(b)); the input alphabet $\Sigma$, used to generate both inputs and string rewrite programs; the minimum and maximum input string lengths $l_{min}$ and $l_{max}$ ($l_{min}=2$, $l_{max}=6$ in Fig 1(b)); the minimum and maximum cascade lengths $L_{min}$ and $L_{max}$ ($L_{min}\leq|\vec{p}|\leq L_{max}$; $|\vec{p}|=4$, $L_{min}=2$, $L_{max}=5$ in Fig 1(b)); the minimum and maximum substring lengths $s_{min}$ and $s_{max}$ for $\alpha$ and $\beta$ in the rewrite programs ($s_{min}\leq|\alpha|,|\beta|\leq s_{max}$); and the desired dataset size $D$. We denote a benchmark snapshot generated with fixed parameter values as $\mathcal{D}(n,\Sigma,L_{min},L_{max},l_{min},l_{max},s_{min},s_{max},D)$.
For the program reordering task, we take PBE instances (Fig 1(a), second from top) generated by the proposer and permute the program sequence by swapping program pairs that exhibit feeding or bleeding interactions. This produces a permuted sequence $\sigma(\vec{p})$ and a different output vector $\vec{o}_{\sigma}=\sigma(\vec{p})(\vec{\imath})\neq\vec{o}$. The goal of the LLM is to recover the original program sequence $\vec{p}$ from the original inputs $\vec{\imath}$, original outputs $\vec{o}$, and the permuted program sequence $\sigma(\vec{p})$, such that $\vec{p}(\vec{\imath})=\vec{o}$.
@@ 3.1.1 Input Sampling
Our input sampling procedure is parameterized by $(n,\Sigma,l_{min},l_{max})$ as defined in Section 3.1. We build the initial set of input strings $\vec{i_{0}}=\langle i^{0}_{1},\dots,i^{0}_{n}\rangle$ ($hwwgdb,jh,\dots wcw$ in Fig 1(b)) by independently sampling each input $i^{0}_{j}$ for $1\leq j\leq n$. For each $i^{0}_{j}$, we first sample its length from a uniform discrete distribution, $|i^{0}_{j}|\sim\mathrm{Unif}\{l_{min},\dots,l_{max}\}$, where $|i^{0}_{j}|$ denotes the length of the string $i^{0}_{j}$. We then generate the string itself by sampling $|i^{0}_{j}|$ characters independently with replacement from the alphabet $\Sigma$, i.e., $i^{0}_{j}\sim\mathrm{Unif}(\Sigma)^{|i^{0}_{j}|}$.
@@ 3.1.2 Program Sampling and Output Generation
Program sampling is parameterized by $(\Sigma,L_{min},L_{max},s_{min},s_{max})$. We begin by programmatically selecting a cascade length $L$ such that $L_{min}\leq L\leq L_{max}$ ($L=4$ in Fig 1(b)). Next, we construct a sequence of $L$ programs $\vec{p}=\langle p_{1},\dots,p_{L}\rangle$, where each program is of the form $p_{k}=\mathrm{replace}(\alpha_{k},\beta_{k})$. Here, $\mathrm{replace}$ has the same semantics as Python’s built-in replace() method for strings: the substring $\alpha_{k}$ is replaced by $\beta_{k}$, with the restriction that $\alpha_{k}$ is non-empty. For example in Fig 1(b), the first program $c\rightarrow wa$ is parameterized by $\alpha_{1}=c$ and $\beta_{1}=wa$. To sample each program $p_{k}$, we generate $\alpha_{k}$ and $\beta_{k}$ independently, following a procedure analogous to input sampling:
@@
    Sample $\alpha_{k}$ uniformly from the set of substrings of length $|\alpha_{k}|$ in the intermediate input vector $\vec{\imath}_{k-1}=p_{k-1}(\dots p_{1}(\vec{\imath}))$ (where $\mathrm{Substr}_{s}(\vec{\imath}_{k-1})$ denotes the set of all substrings of length $s$ in $\vec{\imath}_{k-1}$): $\alpha_{k}\sim\mathrm{Unif}(\mathrm{Substr}_{|\alpha_{k}|}(\vec{\imath}_{k-1}))$ (note that $\vec{\imath}_{0}=\vec{\imath}$). For example in Fig 1(b), for $\alpha_{2}$, we had $|\alpha_{2}|=1$, thus we would sample uniformly from substrings of length one (basically all characters) present in the intermediate inputs $\vec{i_{1}}=\langle hwwgdb,jh,ahwawb,xuuiib,wwaw\rangle$.
@@
The outputs are then generated by executing the program cascade $\vec{p}$ over the inputs $\vec{\imath}$ as $\vec{o}=\vec{p}(\vec{\imath})$ which as stated above is just recursive application of the replace programs $p_{k}(p_{k-1}(\dots p_{1}(\vec{i_{o}})))$